OpenStudy (cutiecomittee123):

P(h)=P*oe^-0.00012h In the function Po is the air pressure at the surface of the earth and h is the height above the surface of the earth, measured in meters. At what height will the air pressure be 50% of the air pressure on the earths surface? 4166.7m 0.59m 2148.9m 5776.2m

OpenStudy (irishboy123):

\[P(h)=P_o e^{-0.00012h}\] did you mean this? if so, do you see the solution?!

OpenStudy (cutiecomittee123):

yes that is what i meant

OpenStudy (cutiecomittee123):

so isnt it just P9h)=.50^-0,00012h

OpenStudy (cutiecomittee123):

P(h)=.50^-0.00012h

OpenStudy (cutiecomittee123):

@IrishBoy123 now what?

OpenStudy (irishboy123):

\[P(h) = P_{o} e^{−0.00012 \ h} \\ P(0)=P_{o} e^{−0.00012 \ (0)} = Po \\ P(H)=P_{o} e^{−0.00012 \ H} = \frac {1}{2}Po \] \[e^{−0.00012 \ H} = \frac {1}{2}\] Solve

OpenStudy (cutiecomittee123):

so ln(.5)/-0.00012

OpenStudy (irishboy123):

-0.00012H = ln (1/2) 0.00012H = ln (2) H = ln (2) / 0.00012 same thing as yours but do you really agree with this? it's your project after all!!

OpenStudy (cutiecomittee123):

yes i agree